OpenStudy (anonymous):

use laplace transformation to solve the given initial value problem y''-2y'+2y=e^-t

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):

whats the initial value ?

OpenStudy (dan815):

u take laplace transform of the left and right

OpenStudy (dan815):

and u get to solve the f(t) given the inverse laplace transform

OpenStudy (dan815):

L(f')=sL(F)-f(0) L(f'')=sL(f')-f'(0) =s*(sL(F)-f(0))-f'(o)

OpenStudy (dan815):

ok i derived another equ

OpenStudy (dan815):

L(y'')=s^2L(y) -sy(0)-y'(0)

OpenStudy (dan815):

L{y^n}=s^nL(y)-s^(n-1)y(0)-s^(n-2)*y'(0).... and so on

OpenStudy (anonymous):

initial values are y(0)=0 y'(0)=1 sorry for the delay

OpenStudy (dan815):

if n is the nth differential of y

OpenStudy (dan815):

its ok dont let it happen again

OpenStudy (dan815):

y''-2y'+2y=e^-t s^2Y-sy(0)-y'(0)-2sY-y(0)=1/(s+1) (s^2-2s)*Y-s*0-1-0=1/(s+1)

OpenStudy (dan815):

Y=(1/(s+1) +1) * 1/(s^2-2s)

OpenStudy (dan815):

solve or Y

OpenStudy (dan815):

and then inverse laplace t

OpenStudy (dan815):

make Y look nicer so u can apply transform

OpenStudy (dan815):

notice the capital Y and small y , Y is the L{y}

OpenStudy (anonymous):

for start i wrote s^2Y(s)-sY(0)-Y'(0)-2[sY(s)-y(0)]+2Y(s)=1/s+1

OpenStudy (dan815):

oh theres a 2 ys oops :)

OpenStudy (dan815):

okay so all that in

OpenStudy (anonymous):

yeap you missed that ;p

OpenStudy (anonymous):

i don't know you to continue

OpenStudy (dan815):

(s^2-2s+2)Y-1=1/(s+1)

OpenStudy (anonymous):

agree

OpenStudy (dan815):

okay so

OpenStudy (dan815):

Y=?

OpenStudy (dan815):

u will neeed to do partial practions

OpenStudy (dan815):

that bracket is not nice too umm

OpenStudy (dan815):

u will hav to make its a perfect square form then

OpenStudy (dan815):

and apply the shift formula

OpenStudy (anonymous):

yes that not good that's why i could not finish it

OpenStudy (dan815):

(s^2-2s+2)Y=1/(s+1) +1 ((s-1)^2+2-1)Y=1/(s+1) +1

OpenStudy (dan815):

now its a nice form

OpenStudy (anonymous):

yes now it's better

OpenStudy (dan815):

where s-1 is the variable basically

OpenStudy (dan815):

or in other words e^at shift factor

OpenStudy (dan815):

now do partial fraction decomposistion

OpenStudy (dan815):

((s-1)^2+2-1)Y=1/(s+1) +1 Y=(1/(s+1) +1)*1/((s-1)^2+1)

OpenStudy (dan815):

rewrite 1/((s+1)((s-1)^2+1)))=A/(s+1)+(Bx+C)/(s-1)^2+1)

OpenStudy (dan815):

and u will have a 1/((s-1)^2+1) byitself too

OpenStudy (dan815):

that u already knw just see how the Bx+c part works out

OpenStudy (anonymous):

ok thank you very much

OpenStudy (anonymous):

so i will have 1=A(s-1)^2+(BX+C(s+1)+(s-1)^2?

OpenStudy (dan815):